# A Comprehensive Survey for Hyperspectral Image Classification: The Evolution from Conventional to Transformers

### v3

@@ 12 Explainable AI and Interoperability
Explainable AI (EAI) and interpretability play crucial roles in HSC, enabling users to understand and trust the decision-making process of the models [369, 370, 371, 372, 373, 374, 375, 376, 377, 378]. Here is a detailed discussion on EAI and interpretability in this context:
@@
    Interpretability Techniques: Various techniques are employed to achieve interpretability in HSC [379, 380, 381, 51, 382]:
@@
      Feature Importance: Feature importance methods, such as permutation importance [383] or feature saliency maps [384], help identify the most influential spectral bands or spatial locations contributing to the classification decision. These methods provide insights into which features are critical for classification and can guide further analysis or data collection.
@@
      Attention Mechanisms: Attention mechanisms [385, 386, 68], commonly used in Transformer-based models, allow users to visualize the attention weights assigned to different spectral bands or spatial locations. By highlighting the relevant regions, attention maps provide interpretability by revealing which parts of the HSI are crucial for classification.
@@
      Rule Extraction: Rule extraction techniques aim to extract human-understandable rules or decision trees from complex models [387]. These rules provide explicit conditions that determine the classification outcome for a given input, making the decision process more transparent and interpretable.
@@
    Domain Expert Collaboration: In HSC, the collaboration between AI experts and domain experts is crucial to achieving interpretability [388, 389]. Domain experts possess in-depth knowledge of the spectral signatures and spatial characteristics of the objects or materials of interest. By involving domain experts in the interpretability process, their expertise can be leveraged to validate and refine the explanations provided by the AI models, ensuring that the interpretations align with domain-specific knowledge.
@@
    Transparency and Trust: EAI and interpretability techniques help build transparency and trust in HSC models [273, 390, 391, 392]. By providing understandable and justifiable explanations, users can assess the model’s reliability, understand its limitations, and make informed decisions based on the model’s outputs. This transparency is particularly important in critical applications, such as environmental monitoring, precision agriculture, or target detection, where the consequences of misclassifications can be significant.
@@
    Regulatory and Ethical Considerations: In some domains, regulatory frameworks or ethical considerations may require explanations for AI-based decisions [393]. For example, in healthcare or remote sensing applications, it may be necessary to explain the reasons behind a diagnosis or land cover classification. Incorporating interpretability techniques in HSC models can help meet requirements and ensure compliance with regulations and ethical standards by providing insights into how the model arrives at its decisions. Interpretability techniques offer transparency by explaining which features or spectral signatures are influential in the classification process. This transparency enables users to understand the model’s reasoning, identify potential biases or errors, and ensure that decisions align with ethical and regulatory guidelines. Additionally, interpretability techniques facilitate model validation and verification, aiding in the assessment of model robustness and reliability.
@@
    Trade-off with Performance: Achieving high interpretability may come with a trade-off in terms of model performance [394]. Simpler models or interpretable methods might not capture the complexity of HS data as effectively as more complex models.
@@
    Interpretable DL: DL models, particularly CNNs and Transformers, can be challenging to interpret due to their hierarchical and non-linear nature [379, 395, 396]. Balancing interpretability with the power of DL is an ongoing research challenge.
@@
    Human Understanding: The interpretability provided by a model should align with the human understanding of the data [374]. It is essential to present information in a way that is meaningful and relevant to domain experts and end-users.
@@ 13 Diffusion Models for Hyperspectral Imaging
Diffusion models for HS imaging are a class of computational techniques used to enhance and analyze HSIs [397, 398, 399, 400, 401, 402, 403]. Diffusion models aim to exploit the underlying spatial and spectral information present in HS data to improve image quality, denoise noisy images, and enhance valuable spectral features for various applications such as remote sensing, environmental monitoring, and object detection. Diffusion models are typically based on the principles of diffusion processes, which simulate the spreading and smoothing of information within an image. These models leverage the notion that neighboring pixels in an image are likely to have similar spectral characteristics. By propagating information across neighboring pixels, diffusion models can effectively enhance the spatial and spectral coherence of HSIs.
@@
One popular diffusion model used in HS imaging is the Perona-Malik (PM) diffusion model [404, 405, 406]. The PM model is based on the heat equation and aims to remove noise while preserving meaningful image structures. It achieves this by diffusing image gradients according to a diffusion coefficient that is adaptively modulated by the local image characteristics. The PM model has been successfully applied to HS denoising tasks, effectively reducing noise while preserving important spectral features.
@@
Another widely used diffusion model is the Total Variation (TV) model [407, 408, 401, 409, 410]. The TV model seeks to denoise HSIs by minimizing the TV of the image, which measures the sum of the absolute differences between neighboring pixel values. By promoting piecewise smoothness in the image, the TV model can effectively suppress noise while preserving sharp edges and fine details.
@@
In addition to denoising, diffusion models can also be employed for other tasks in HS imaging, such as image fusion and feature extraction [399, 411, 412, 413]. For image fusion, diffusion-based methods can be used to integrate information from multiple HSIs or different modalities, thereby enhancing the overall image quality and reducing noise [414, 406]. In terms of feature extraction, diffusion models can be employed to enhance discriminative spectral features, making them more distinguishable for subsequent classification or target detection tasks [415, 416, 417, 414, 418].
@@
It is worth noting that while diffusion models have shown promising results in HS imaging, they also come with certain limitations. These models may struggle with preserving fine details and subtle spectral variations, as excessive diffusion can lead to over-smoothing [419]. Additionally, diffusion models can be computationally intensive, especially when applied to large-scale HS datasets, requiring efficient algorithms and parallel computing techniques to ensure real-time or near-real-time processing. In summary, diffusion models provide valuable tools for enhancing HSIs by leveraging spatial and spectral information. These models, such as the PM and TV models, offer effective denoising, image fusion, and feature extraction capabilities. However, careful parameter tuning and algorithm optimization are necessary to strike a balance between noise suppression and preservation of important spectral details. Ongoing research continues to explore and refine diffusion models for HS imaging to address their limitations and expand their applicability in various domains.
@@ 14 Open Challenges and Research Questions
    Limited Labeled Training Data: Challenge: Acquiring labeled training data for HSIs can be expensive and time-consuming. Limited labeled data may hinder the development of accurate and robust classification models [420, 91]. Research Questions: How can transfer learning [421] be effectively applied to leverage knowledge from related domains or existing labeled datasets to improve classification performance with limited labeled HS data? What AL strategies can be employed to strategically select and query the most informative samples, optimizing the use of limited labeled data?
@@
    Curse of Dimensionality: Challenge: HS data typically exhibits high dimensionality due to the large number of spectral bands. This high-dimensional space can lead to the curse of dimensionality, affecting the performance of traditional classification algorithms [422]. Research Questions: How can dimensionality reduction techniques be tailored to HS data to preserve relevant information while mitigating the challenges associated with the curse of dimensionality? What DL architectures or AE can effectively capture and represent the intrinsic structure of HS data in a lower-dimensional space?
@@
    Spectral-Spatial Complexity: Challenge: HSIs often contain intricate spectral-spatial patterns that are challenging to capture using traditional classifiers [181, 423]. Research Questions: How can advanced machine learning models, such as DNNs, be designed to integrate spectral and spatial information for improved classification accuracy effectively? What role do graph-based models or attention mechanisms play in capturing spectral-spatial dependencies, and how can they be optimized for HSC?
@@
    Class Imbalance and Rare Events: Challenge: Class imbalance is prevalent in HS datasets, where certain classes may have fewer instances. Additionally, rare events or anomalies may be crucial but challenging to detect [424, 425]. Research Questions: How can HSI classifiers handle class imbalance, and what techniques can be employed to address the challenge of rare events or anomalies in the data? What role can generative models, such as GANs, play in augmenting rare class samples and enhancing classification performance?
@@
    Robustness to Environmental Variability: Challenge: HS data may be sensitive to variations in environmental conditions, such as lighting, atmospheric effects, and seasonal changes, leading to reduced model generalization [426, 427]. Research Questions: How can HS classifiers be made robust to environmental variability, and what preprocessing or normalization techniques are effective in reducing the impact of these variations? What transfer learning strategies can be employed to adapt models trained in one environment to perform well in different environmental conditions?
@@
    Explainability and Interpretability: Challenge: DL models and complex classifiers may lack interpretability, hindering user understanding and trust in the classification results [379, 380, 381, 51, 382]. Research Questions: What methods and tools can be developed to provide explainability and interpretability in HSC, especially for DL models? How can the interpretability of models be aligned with the specific needs and expectations of domain experts and end-users in different application domains?
@@
    Real-Time Processing and Resource Constraints: Challenge: In certain applications, HSC needs to operate in real-time or under resource constraints, necessitating efficient algorithms [428, 429, 430, 431]. Research Questions: How can lightweight and efficient models be designed for real-time HSC, considering the constraints of processing power, memory, and energy consumption? What trade-offs can be made between model complexity and accuracy to achieve optimal performance in resource-constrained environments?

### v4

@@ 12 Explainable AI and Interoperability
Explainable AI (EAI) and interpretability play crucial roles in HSC, enabling users to understand and trust the decision-making process of the models (Turan et al.,, 2023; Zhang and Abdulla,, 2023; Hong et al., 2024a, ). Here is a detailed discussion on EAI and interpretability in this context:
@@
    Interpretability Techniques: Various techniques are employed to achieve interpretability in HSC (Gong et al.,, 2024; Deshpande et al.,, 2024):
@@
      Feature Importance: Feature importance methods, such as permutation importance (Putkiranta et al.,, 2024) or feature saliency maps (Zhou et al., 2024b, ), help identify the most influential spectral bands or spatial locations contributing to the classification decision. These methods provide insights into which features are critical for classification and can guide further analysis or data collection.
@@
      Attention Mechanisms: Attention mechanisms (Wang et al., 2024d, ; Yang et al.,, 2024), commonly used in Transformer-based models, allow users to visualize the attention weights assigned to different spectral bands or spatial locations. By highlighting the relevant regions, attention maps provide interpretability by revealing which parts of the HSI are crucial for classification.
@@
      Rule Extraction: Rule extraction techniques aim to extract human-understandable rules or decision trees from complex models (Castano et al.,, 2024). These rules provide explicit conditions that determine the classification outcome for a given input, making the decision process more transparent and interpretable.
@@
    Domain Expert Collaboration: In HSC, the collaboration between AI experts and domain experts is crucial to achieving interpretability (Zhang et al., 2024a, ; Yu et al., 2024a, ). Domain experts possess in-depth knowledge of the spectral signatures and spatial characteristics of the objects or materials of interest. By involving domain experts in the interpretability process, their expertise can be leveraged to validate and refine the explanations provided by the AI models, ensuring that the interpretations align with domain-specific knowledge.
@@
    Transparency and Trust: EAI and interpretability techniques help build transparency and trust in HSC models (Rubab et al.,, 2024; Tejasree and Agilandeeswari,, 2024; Panda et al.,, 2024). By providing understandable and justifiable explanations, users can assess the model’s reliability, understand its limitations, and make informed decisions based on the model’s outputs. This transparency is particularly important in critical applications, such as environmental monitoring, precision agriculture, or target detection, where the consequences of misclassifications can be significant.
@@
    Regulatory and Ethical Considerations: In some domains, regulatory frameworks or ethical considerations may require explanations for AI-based decisions (Ayala et al.,, 2023). For example, in healthcare or remote sensing applications, it may be necessary to explain the reasons behind a diagnosis or land cover classification. Incorporating interpretability techniques in HSC models can help meet requirements and ensure compliance with regulations and ethical standards by providing insights into how the model arrives at its decisions. Interpretability techniques offer transparency by explaining which features or spectral signatures are influential in the classification process. This transparency enables users to understand the model’s reasoning, identify potential biases or errors, and ensure that decisions align with ethical and regulatory guidelines. Additionally, interpretability techniques facilitate model validation and verification, aiding in the assessment of model robustness and reliability.
@@
    Trade-off with Performance: Achieving high interpretability may come with a trade-off in terms of model performance (Hanchate et al.,, 2023). Simpler models or interpretable methods might not capture the complexity of HS data as effectively as more complex models.
@@
    Interpretable DL: DL models, particularly CNNs and Transformers, can be challenging to interpret due to their hierarchical and non-linear nature (Xie et al.,, 2022; Tian et al.,, 2023; Li et al., 2023a, ). Balancing interpretability with the power of DL is an ongoing research challenge.
@@
    Human Understanding: The interpretability provided by a model should align with the human understanding of the data (Oveis et al.,, 2023). It is essential to present information in a way that is meaningful and relevant to domain experts and end-users.
@@ 13 Diffusion Models for Hyperspectral Imaging
Diffusion models for HS imaging are a class of computational techniques used to enhance and analyze HSIs (Miao et al.,, 2023; Wu et al.,, 2023; Chen et al., 2023b, ; Liu et al., 2023b, ; Murphy and Maggioni,, 2019; Cao et al., 2024c, ; Zhang et al., 2024b, ). Diffusion models aim to exploit the underlying spatial and spectral information present in HS data to improve image quality, denoise noisy images, and enhance valuable spectral features for various applications such as remote sensing, environmental monitoring, and object detection. Diffusion models are typically based on the principles of diffusion processes, which simulate the spreading and smoothing of information within an image. These models leverage the notion that neighboring pixels in an image are likely to have similar spectral characteristics. By propagating information across neighboring pixels, diffusion models can effectively enhance the spatial and spectral coherence of HSIs.
@@
One popular diffusion model used in HS imaging is the Perona-Malik (PM) diffusion model (Wen et al.,, 2020; Sayah et al.,, 2021; Wang et al.,, 2010). The PM model is based on the heat equation and aims to remove noise while preserving meaningful image structures. It achieves this by diffusing image gradients according to a diffusion coefficient that is adaptively modulated by the local image characteristics. The PM model has been successfully applied to HS denoising tasks, effectively reducing noise while preserving important spectral features.
@@
Another widely used diffusion model is the Total Variation (TV) model (Duan et al.,, 2019; Peng et al.,, 2022; Murphy and Maggioni,, 2019; Liu et al., 2019b, ; Wang et al.,, 2018). The TV model seeks to denoise HSIs by minimizing the TV of the image, which measures the sum of the absolute differences between neighboring pixel values. By promoting piecewise smoothness in the image, the TV model can effectively suppress noise while preserving sharp edges and fine details.
@@
In addition to denoising, diffusion models can also be employed for other tasks in HS imaging, such as image fusion and feature extraction (Chen et al., 2023b, ; Chen et al., 2023a, ; Cui et al.,, 2024; Zhou et al., 2024a, ). For image fusion, diffusion-based methods can be used to integrate information from multiple HSIs or different modalities, thereby enhancing the overall image quality and reducing noise (Li et al., 2023b, ; Wang et al.,, 2010). In terms of feature extraction, diffusion models can be employed to enhance discriminative spectral features, making them more distinguishable for subsequent classification or target detection tasks (Wu et al., 2019b, ; Tan et al.,, 2024; Tasissa et al.,, 2021; Li et al., 2023b, ; Wu et al., 2022b, ).
@@
It is worth noting that while diffusion models have shown promising results in HS imaging, they also come with certain limitations. These models may struggle with preserving fine details and subtle spectral variations, as excessive diffusion can lead to over-smoothing (Jeong and Shin,, 2023). Additionally, diffusion models can be computationally intensive, especially when applied to large-scale HS datasets, requiring efficient algorithms and parallel computing techniques to ensure real-time or near-real-time processing. In summary, diffusion models provide valuable tools for enhancing HSIs by leveraging spatial and spectral information. These models, such as the PM and TV models, offer effective denoising, image fusion, and feature extraction capabilities. However, careful parameter tuning and algorithm optimization are necessary to strike a balance between noise suppression and preservation of important spectral details. Ongoing research continues to explore and refine diffusion models for HS imaging to address their limitations and expand their applicability in various domains.
@@
14 State Space Model (Mamba)
The Mamba model represents an innovative approach to HSC, leveraging the state space model (SSM) framework to process spectral and spatial information efficiently. Traditional DL models, such as CNNs and Transformers, have demonstrated strong performance in HSC by capturing spatial and spectral dependencies (Ahmad et al., 2024b, ). However, these models often face challenges with computational complexity and scalability, especially when handling high-dimensional data typical of HSIs. The Mamba model addresses these limitations by utilizing the efficiency of state space operations, offering a powerful alternative to conventional methods.
The Mamba model is fundamentally an SSM adapted for HS data analysis. SSMs are known for their capacity to represent sequential data efficiently by capturing both spatial and temporal dependencies through a set of hidden states. In the context of HSIs, each pixel is represented as a sequence of spectral bands, and Mamba uses state space equations to model this sequence dynamically. The Mamba model is structured with key components: Spectral-Spatial Token Generation: Generates tokens representing both spatial and spectral features, offering a more compact representation of HS data. State Transition Mechanism: Efficiently encodes dependencies among spectral bands, capturing relationships with fewer computations than standard attention mechanisms used in Transformers. Output Layer: Provides a classification output by processing the state representations, which contain sufficient spectral-spatial information for accurate classification. The core of Mamba relies on state space equations, which are expressed as:
$x_{t+1}=Ax_{t}+Bu_{t}$ (2)
$y_{t}=Cx_{t}+Du_{t}$ (3)
where $x_{t}$ represents the state vector at time $t$, $u_{t}$ is the input vector (e.g., spectral feature), $A$ and $B$ are the transition matrices that determine state evolution, and $C$ and $D$ project the state to the output space. This representation enables Mamba to maintain a compact set of states for each spectral band, reducing the need for extensive parameters and computations, as opposed to the dense self-attention layers in Transformer-based models.
The Mamba model offers several significant advantages over Transformers and other conventional models: Reduced Computational Complexity: The Mamba model’s use of state space representations inherently reduces computational requirements. Unlike Transformers, which rely on quadratic complexity in terms of sequence length due to self-attention, Mamba processes sequential data linearly. This is particularly beneficial for HSIs, which have long spectral sequences, making Mamba a faster and more resource-efficient choice. Memory Efficiency: Transformers often require extensive memory for handling large spectral sequences due to the self-attention mechanism’s need to store intermediate representations. The Mamba model, by contrast, uses fewer parameters and requires less memory, enabling it to handle larger HS datasets without sacrificing performance. Ability to Capture Long-Range Dependencies: While CNNs typically struggle with capturing long-range dependencies, the state space formulation in Mamba naturally encodes these dependencies through state transitions, allowing for effective learning of spectral-spatial relationships in HSIs. Improved Scalability: The model’s reduced complexity allows it to scale well with high-resolution HS datasets, which are often challenging for other models. Mamba’s efficient structure enables it to perform well even as the dimensionality of the input data grows, a crucial advantage when dealing with real-world HS data. Noise Robustness: HS data often contain noise, which can impair classification accuracy. The state transitions in Mamba smooth out such noise by maintaining continuity between states. This robustness is further enhanced by the token generation process, which aggregates spectral and spatial information, providing resilience to small variations in the input.
The MHSSMamba (Ahmad et al., 2024b, ) model enhances Spatial-Spectral Mamba by incorporating multi-head self-attention and token enhancement, effectively capturing complex spectral-spatial relationships and managing long-range dependencies in hyperspectral data. This approach improves computational efficiency while preserving critical contextual information across spectral bands. WaveMamba (Ahmad et al., 2024f, ) introduces an innovative integration of wavelet transformation with Spatial-Spectral Mamba architecture, enhancing hyperspectral image classification by capturing both local textures and global contexts. The Spatial-Spectral Morphological Mamba (MorpMamba) (Ahmad et al., 2024c, ) model optimizes HSC by integrating morphological operations and an SSM for computational efficiency, addressing the limitations of Transformers. Through spatial-spectral token generation, feature enhancement, and multi-head self-attention, MorpMamba effectively fuses detailed structural and spectral information, improving classification and producing high-quality ground truth maps.
The MambaHSI (Li et al., 2024i, ) model leverages Mamba’s efficient long-range modeling capabilities, introducing spatial and spectral Mamba blocks to adaptively capture HSI data’s spatial-spectral interactions. This approach, validated across four HSI datasets, demonstrates MambaHSI’s superiority in classification accuracy and highlights Mamba’s potential as a next-generation backbone for HSI models. The Mamba-in-Mamba (MiM) (Zhou et al.,, 2025) architecture pioneers a new approach for HSC, utilizing an efficient Mamba-Cross-Scan (MCS) transformation and tokenized feature learning to overcome computational challenges. By integrating multi-scale loss and enhanced semantic tokenization, MiM offers robust performance, especially suited for remote sensing tasks requiring high precision with limited training data.
The HyperMamba model (Liu et al.,, 2024) introduces a spectral-spatial adaptive Mamba for HSC, balancing spatial and spectral feature extraction efficiently. By employing Spatial Neighborhood Adaptive Scanning (SNAS) and Spectral Adaptive Enhancement Scanning (SAES), HyperMamba effectively captures intricate spatial-spectral dependencies, achieving state-of-the-art performance across multiple HSI datasets. The Local Enhanced Mamba (LE-Mamba) network (Wang et al., 2024a, ) offers an efficient solution for HSC by combining a Local Enhanced Spatial SSM (LES-S6) and Central Region Spectral SSM (CRS-S6) for precise spatial-spectral integration, while the Multi-Scale Convolutional Gated Unit (MSCGU) ensures robust feature aggregation, achieving strong non-causal feature extraction and enhanced classification performance.
The HLMamba (Liao et al.,, 2024) model leverages a gradient joint algorithm (GJA) for edge contour extraction and a multimodal Mamba fusion module (MMFM) for efficient feature integration across HSI and LiDAR data, enhancing classification performance by capturing complex interrelationships and long-distance dependencies in multimodal data. The spatial-spectral interaction super-resolution CNN-Mamba fusion network (Zhao et al., 2024a, ) enhances the resolution of low spatial resolution HSIs and high spatial resolution multispectral images through mutual guidance, ensuring effective computational efficiency in extracting both local and global features for improved image fusion.
In HSC, computational efficiency is crucial due to the high dimensionality of the data. The Mamba model demonstrates not only comparable or superior classification accuracy to Transformers and CNNs but also achieves this with lower computational cost. This makes Mamba a practical solution for resource-constrained applications, such as onboard processing in drones or real-time monitoring systems. Additionally, due to its compact state space representation, the model can be implemented in edge devices for field applications in agriculture, environmental monitoring, and disaster response, where power efficiency and computational efficiency are essential.
14.a Current Limitations of the Mamba Model:
Scalability to High Dimensions: While the Mamba model exhibits linear complexity compared to traditional Transformers, it may still face challenges when scaling to extremely high-dimensional HS data. As the number of spectral bands increases, the model’s performance can degrade due to the increased complexity of capturing the intricate spectral relationships.
Integration of Temporal Dynamics: The original Mamba model was primarily designed for spatial and spectral interactions, which limits its capability to effectively incorporate temporal dynamics in HSI sequences. This is particularly relevant in applications involving dynamic changes in the environment, such as agriculture or urban development.
Limited Robustness to Noisy Data: Although the Mamba architecture demonstrates strong performance under controlled conditions, it may not be robust enough to handle noisy or corrupted HS data, which is common in real-world scenarios. This limitation can lead to suboptimal classification outcomes, especially in the presence of atmospheric disturbances or sensor artifacts.
Feature Representation and Interpretation: The model’s representation of features, particularly in terms of interpretability, can be limited. While Mamba models capture complex relationships, understanding how these relationships contribute to final decisions remains a challenge, especially in critical applications where interpretability is essential.
Heterogeneous Data Fusion: Current Mamba implementations may struggle with the effective fusion of multimodal data (e.g., combining HSIs with LiDAR or RGB data). The model’s architecture needs to be adapted to seamlessly integrate heterogeneous features while preserving important spatial and spectral information.
14.b Possible Future Research Directions
Enhanced Feature Representation Techniques: Future research could focus on developing advanced feature extraction methods that explicitly address the high-dimensional nature of HS data. This could include the integration of hybrid architectures to enhance the model’s ability to extract meaningful features.
Temporal Mamba Extensions: To address the limitations regarding temporal dynamics, researchers could explore the incorporation of temporal models into the Mamba framework. This could involve developing a spatiotemporal Mamba variant that effectively models changes over time, enabling applications in dynamic monitoring and analysis of environments.
Robustness to Noisy Data: Investigating noise-resistant training strategies, such as adversarial training or data augmentation techniques, can enhance the robustness of the Mamba model. Research could focus on developing specific regularization techniques tailored for HS data that improve classification performance in noisy conditions.
Interpretable Mamba Architectures: Efforts should be made to design interpretable Mamba architectures that can elucidate the reasoning behind classification decisions. Techniques such as attention visualization, feature importance mapping, and explainable AI frameworks can be integrated to improve transparency in model predictions.
Multimodal and Heterogeneous Data Fusion: Future work could investigate the extension of Mamba models to handle the integration of various types of data more effectively. This could involve developing new fusion strategies that can dynamically adapt to different data sources while maintaining the spatial-spectral integrity of HSIs.
@@ 15 Open Challenges and Research Questions
    Limited Labeled Training Data: Challenge: Acquiring labeled training data for HSIs can be expensive and time-consuming. Limited labeled data may hinder the development of accurate and robust classification models (Xiao et al.,, 2024; Ye et al.,, 2024). Research Questions: How can transfer learning (Rise et al.,, 2024) be effectively applied to leverage knowledge from related domains or existing labeled datasets to improve classification performance with limited labeled HS data? What AL strategies can be employed to strategically select and query the most informative samples, optimizing the use of limited labeled data?
@@
    Curse of Dimensionality: Challenge: HS data typically exhibits high dimensionality due to the large number of spectral bands. This high-dimensional space can lead to the curse of dimensionality, affecting the performance of traditional classification algorithms (Hong et al., 2019a, ). Research Questions: How can dimensionality reduction techniques be tailored to HS data to preserve relevant information while mitigating the challenges associated with the curse of dimensionality? What DL architectures or AE can effectively capture and represent the intrinsic structure of HS data in a lower-dimensional space?
@@
    Spectral-Spatial Complexity: Challenge: HSIs often contain intricate spectral-spatial patterns that are challenging to capture using traditional classifiers (Zhao et al., 2024b, ; Song et al., 2024a, ). Research Questions: How can advanced machine learning models, such as DNNs, be designed to integrate spectral and spatial information for improved classification accuracy effectively? What role do graph-based models or attention mechanisms play in capturing spectral-spatial dependencies, and how can they be optimized for HSC?
@@
    Class Imbalance and Rare Events: Challenge: Class imbalance is prevalent in HS datasets, where certain classes may have fewer instances. Additionally, rare events or anomalies may be crucial but challenging to detect (Chang et al.,, 2024; Hu et al.,, 2024). Research Questions: How can HSI classifiers handle class imbalance, and what techniques can be employed to address the challenge of rare events or anomalies in the data? What role can generative models, such as GANs, play in augmenting rare class samples and enhancing classification performance?
@@
    Robustness to Environmental Variability: Challenge: HS data may be sensitive to variations in environmental conditions, such as lighting, atmospheric effects, and seasonal changes, leading to reduced model generalization (Yu and Cui,, 2024). Research Questions: How can HS classifiers be made robust to environmental variability, and what preprocessing or normalization techniques are effective in reducing the impact of these variations? What transfer learning strategies can be employed to adapt models trained in one environment to perform well in different environmental conditions?
@@
    Explainability and Interpretability: Challenge: DL models and complex classifiers may lack interpretability, hindering user understanding and trust in the classification results (Gong et al.,, 2024; Deshpande et al.,, 2024). Research Questions: What methods and tools can be developed to provide explainability and interpretability in HSC, especially for DL models? How can the interpretability of models be aligned with the specific needs and expectations of domain experts and end-users in different application domains?
@@
    Real-Time Processing and Resource Constraints: Challenge: In certain applications, HSC needs to operate in real-time or under resource constraints, necessitating efficient algorithms (Hsu et al.,, 2024; Li et al., 2024c, ; Zhang et al., 2024c, ; Martins et al.,, 2024). Research Questions: How can lightweight and efficient models be designed for real-time HSC, considering the constraints of processing power, memory, and energy consumption? What trade-offs can be made between model complexity and accuracy to achieve optimal performance in resource-constrained environments?